OpenStudy (anonymous):

Use the fundamental identities to simplify the expression. There is more than one correct form of the answer. 9sec^2 x (1−sin^2x)

OpenStudy (anonymous):

start with the fact that \[\sin^2(x)+\cos^2(x)=1\] or \[\cos^2(x)=1-\sin^2(x)\] as your first replacement

OpenStudy (anonymous):

get \[9\sec^2(x)\cos^2(x)\] one more step and you are done

OpenStudy (anonymous):

just that the 9 part confused me

OpenStudy (anonymous):

ignore it you get \[9\frac{1}{\cos^2(x)}\times \cos^2(x)=9\]